What is the probability of finding a particle at centre of box in it's first excited state in a infinite potential well ??


1
Expert's answer
2021-03-01T11:52:01-0500

The wave function of particle in the first excited state is

"\psi(x)=\sqrt{\frac{2}{L}}sin(\frac{2\pi x}{L})"

The probability of finding a particle at centre of box  is

"P=|\psi(x)|^2(L)"

"=|\sqrt{\frac{2}{L}}sin(\frac{2\pi (x/2)}{L})|^2(L) =0" since "sin(\pi) =0"